Spectrum:

Think of a number, then double it, then multiply by 5, and knock off the last digit. 1-10 only.

Spectrum:

bet its the number you choose to begin with.

TheOfficialDevan12:

nah dip

Spectrum:

3,6,30,3

Spectrum:

4,8,40,4

Icyhot:

7+7=14 14*5=70 70-0=7

ThatWeirdo:

1+1=2x5=10=1 i give up you wizard